Compute the correlation coefficient of the data below.




*number of years vs. selling price



X 1 3 5 7 9



y 27 23 25 20 15





1
Expert's answer
2021-12-23T08:47:54-0500
"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c}
 & X & Y & XY & X^2 & Y^2 \\ \hline
 & 1 & 27 & 27 & 1 & 729 \\
 \hdashline
 & 3 & 23 & 69 & 9 & 529 \\
 \hdashline
 & 5 & 25 & 125 & 25 & 625 \\
 \hdashline
 & 7 & 20 & 140 & 49 & 400 \\
 \hdashline
 & 9 & 15 & 135 & 81 & 225 \\
 \hdashline
Sum= & 25 & 110 & 496 & 165 & 2508 \\
 \hdashline
\end{array}"

"\bar{X}=\dfrac{1}{n}\displaystyle\sum_{i=1}^nX_i=\dfrac{25}{5}=5"

"\bar{Y}=\dfrac{1}{n}\displaystyle\sum_{i=1}^nY_i=\dfrac{110}{5}=22"

"SS_{XX}=\displaystyle\sum_{i=1}^nX_i^2-\dfrac{1}{n}(\displaystyle\sum_{i=1}^nX_i)^2=165-\dfrac{25^2}{5}=40"

"SS_{YY}=\displaystyle\sum_{i=1}^nY_i^2-\dfrac{1}{n}(\displaystyle\sum_{i=1}^nY_i)^2=2508-\dfrac{110^2}{5}=88"

"SS_{XY}=\displaystyle\sum_{i=1}^nX_iY_i-\dfrac{1}{n}(\displaystyle\sum_{i=1}^nX_i)(\displaystyle\sum_{i=1}^nY_i)"

"=496-\dfrac{25(110)}{5}=-54"

Correlation coefficient:


"r=\dfrac{SS_{XY}}{\sqrt{SS_{XX}}\sqrt{SS_{YY}}}=\dfrac{-54}{\sqrt{40}\sqrt{88}}"

"\approx -0.91017"




"-0.7<r<-11.0"

"r=-0.91017," strong negative correlation.